N2 molecule is spherical of radius 100pm a) what is the volume (in cm3 ) of one mole of molecules of gas

Grade:11

1 Answers

Arun
6 years ago
Dear Aditya
N2 is in spherical form, 
Hence
Volume of one atom = (4/3) *\pi * r³
= 4*22*(10-10)³/3*7 = 4.19 * 10-30 m³
Now for one mole, we muktiply by avagadro number Na.
Hence 
total volume of one mole= 6.023 * 10²³ * 4.19* 10-30
 = 25.23 * 10-7 m³
 
Regards
Arun(askIITians forum expert)